# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ 6.1.1 Mathematical Formulation of Synthetic Data Generation
where $\bm{\mu}_{1}=(2,2)^{T}$ is the cluster center and the radius scaling factor is $0.5$.
@@
where $\bm{\mu}_{2}=(8,2)^{T}$, $a_{2}=1.5$ and $b_{2}=0.4$ are the semi-major and semi-minor axes respectively.
@@
where $\bm{\mu}_{3}=(2,8)^{T}$ and the horizontal shift of 0.4 in the inner arc creates the crescent shape.
@@
View 2: Complementary Geometric Patterns
View 2 provides complementary information through different geometric patterns positioned to create view-specific clustering challenges: Cluster 1 (Diamond): Generated using a diamond-shaped polar pattern:
@@
where $\bm{\mu}_{4}^{(2)}=(-2,-2)^{T}$, $\sigma=0.3$ is the scale parameter, and $\eta_{i,x},\eta_{i,y}\sim\mathcal{N}(0,1)$.
@@ 6.1.2 Data Generation Properties and Validation
The synthetic data generation framework ensures several critical properties for rigorous algorithm evaluation:
Sample Count Consistency:
Each cluster in both views contains exactly $N_{c}=2,500$ samples, yielding a total dataset size of $N=10,000$ samples per view. This balanced design eliminates bias towards larger clusters.
Geometric Complexity:
The chosen shapes span a range of geometric complexities:
  • Convex structures: Circle, ellipse, diamond
  • Non-convex structures: Crescent, heart, cross
  • Topological variations: Simple connected (most shapes) vs. multiply connected (ring)
  • Parametric complexity: Linear (cross) vs. highly nonlinear (spiral, heart)
View Complementarity:
The positioning strategy ensures that:
@@
where MI denotes mutual information between cluster assignments, ensuring that views provide complementary rather than redundant information.
Noise Characteristics:
Gaussian noise is added with view-specific parameters:
@@ 6.1.3 Federated Data Partitioning Strategy
This partition ensures that:
  • Client 1 receives 85% of the data (8,500 samples)
  • Client 2 receives 15% of the data (1,500 samples)
  • Both clients have access to all four cluster types
  • The partition reflects realistic federated scenarios with heterogeneous data distributions
This synthetic dataset provides a controlled environment for evaluating multi-view clustering algorithms on geometrically complex data while maintaining known ground truth labels for quantitative performance assessment.
Dataset Statistics and Properties
The synthetic multi-view dataset exhibits the following key statistical properties that make it ideal for rigorous algorithm evaluation:
@@
Controlled Experimental Advantages
The carefully designed synthetic framework offers several experimental advantages:
@@
Validation Framework
To ensure the quality and reliability of the synthetic dataset, we implement a comprehensive validation protocol:
@@ 6.1.4 Medical Federated Scenario: Multi-Hospital Patient Data Analysis
Clinical Motivation
Modern healthcare generates vast amounts of multi-modal patient data across different institutions, creating opportunities for collaborative analysis that could improve patient outcomes through enhanced phenotyping and treatment optimization. However, strict privacy regulations (HIPAA, GDPR) and institutional policies prevent direct data sharing between healthcare providers. Our FedHK-MVFC framework addresses this challenge by enabling collaborative clustering analysis while ensuring that sensitive patient data never leaves the originating institution.
Consider a realistic scenario where multiple hospitals seek to collaboratively identify cardiovascular patient phenotypes for improved treatment protocols. Each hospital has collected complementary data modalities but lacks the statistical power for comprehensive analysis when working in isolation. Through federated clustering, these institutions can leverage their collective knowledge while maintaining full compliance with privacy regulations.
Data Modality Mapping
The synthetic dataset’s geometric patterns visualized in Figure 1 serve as proxies for real medical data characteristics:
View 1 (Physiological Measurements): The four distinct shapes in View 1 represent different patient populations based on physiological measurements:
  • Circular cluster: Represents healthy control patients with normal, tightly distributed physiological parameters
  • Elliptical cluster: Models patients with mild cardiovascular risk showing elongated parameter distributions
  • Crescent cluster: Captures patients with complex, non-linear disease progression patterns
  • Spiral cluster: Represents patients with progressive disease states showing temporal evolution
View 2 (Imaging and Behavioral Data): The complementary shapes in View 2 correspond to imaging and behavioral characteristics:
  • Diamond cluster: Structured imaging features with clear anatomical boundaries
  • Ring cluster: Patients with preserved central function but peripheral abnormalities
  • Cross cluster: Behavioral patterns showing distinct lifestyle factor combinations
  • Heart cluster: Complex cardiac structural abnormalities requiring specialized interpretation
Privacy-Preserving Collaborative Analysis
@@
Clinical Workflow Implementation
@@
Multi-View Medical Data Structure
@@
Privacy and Compliance Requirements
@@
Federated Data Distribution Analysis
@@
FedHK-MVFC Advantages in Medical Applications
@@
This medical scenario demonstrates how the synthetic data experiments (with their controlled geometric complexities) translate to real-world healthcare applications where FedHK-MVFC can enable privacy-preserving collaborative analysis across medical institutions.
6.2 Real-World Case Study: A Benchmark Dataset
The Iris dataset is a classic benchmark in machine learning, consisting of 150 samples of iris flowers with four features: sepal length, sepal width, petal length, and petal width. In this case study, we apply the FedHK-MVFC framework to the Iris dataset, simulating a federated learning environment with two clients. The dataset is partitioned such that each client holds a subset of the data, and we aim to cluster the samples into three species: Iris Setosa, Iris Versicolor, and Iris Virginica. We first create a scenario on Iris dataset where attributes 1 and 3 are used for View 1, and attributes 2 and 4 are used for View 2. Then, we distribute the data points between the two clients while ensuring that both views are represented. Each client has access to all three species, but the distribution of samples is uneven, simulating a realistic federated learning scenario. Client one holds 90 samples, while client two holds 60 samples.
@@ 6.3 Experimental Setup and Evaluation Metrics
6.3.1 Baseline Algorithms
We compare our proposed HK-MVFC and FedHK-MVFC algorithms against state-of-the-art multi-view clustering methods:
  • Multi-View Fuzzy C-Means (MV-FCM): Traditional extension of FCM to multi-view data
  • Co-regularized Multi-View Spectral Clustering (CoMSC): Spectral approach with cross-view regularization
  • Federated Multi-View K-Means (FedMV-KMeans): Federated extension of multi-view k-means
  • Multi-View Deep Clustering (MVDC): Deep learning-based multi-view clustering
6.3.2 Evaluation Metrics
Clustering performance is assessed using multiple complementary metrics:
Accuracy-based Metrics:
Accuracy $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[\hat{y}_{i}=y_{i}]$ (116)
Normalized Mutual Information (NMI) $=\frac{2\cdot\text{MI}(\mathbf{y},\hat{\mathbf{y}})}{\text{H}(\mathbf{y})+\text{H}(\hat{\mathbf{y}})}$ (117)
Adjusted Rand Index (ARI) $=\frac{\text{RI}-\mathbb{E}[\text{RI}]}{\max(\text{RI})-\mathbb{E}[\text{RI}]}$ (118)
Geometric Fidelity Metrics:
Silhouette Coefficient $=\frac{1}{N}\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{\max(a_{i},b_{i})}$ (119)
Calinski-Harabasz Index $=\frac{\text{SSB}/(\text{c}-1)}{\text{SSW}/(\text{N}-\text{c})}$ (120)
Multi-View Specific Metrics:
View Consensus $=\frac{1}{S}\sum_{h=1}^{S}\text{NMI}(\hat{\mathbf{y}},\hat{\mathbf{y}}^{(h)})$ (121)
Cross-View Stability $=1-\frac{1}{S(S-1)}\sum_{h\neq h^{\prime}}\|\mathbf{V}^{(h)}-\mathbf{V}^{(h^{\prime})}\|_{2}$ (122)
6.3.3 Computational Complexity Analysis
The synthetic data generation framework exhibits the following computational complexities:
Data Generation Complexity:
For generating $N$ samples across $S$ views with $C$ clusters:
$\mathcal{O}_{generation}$ $=\mathcal{O}(N\cdot S\cdot C_{shape})$ (123)
$\mathcal{O}_{validation}$ $=\mathcal{O}(N\cdot S)$ (124)
$\mathcal{O}_{total}$ $=\mathcal{O}(N\cdot S\cdot(C_{shape}+1))$ (125)
where $C_{shape}$ represents the computational cost of the most complex shape generation (heart curve: $\mathcal{O}(1)$ per sample).
Federated Partitioning Complexity:
$\mathcal{O}_{partition}=\mathcal{O}(N\log N+P\cdot N/P)=\mathcal{O}(N\log N)$ (126)
where $P$ is the number of federated clients.
6.3.4 Theoretical Properties of the Synthetic Framework
Geometric Diversity Theorem:
Theorem 3 (Geometric Coverage).
Let $\mathcal{G}=\{G_{1},G_{2},\ldots,G_{C}\}$ be the set of geometric shapes in our synthetic framework. The framework provides complete coverage of fundamental topological properties:
$\bigcup_{i=1}^{C}\text{Hull}(G_{i})$ $\supset\{\text{convex, non-convex, simply-connected, multiply-connected}\}$ (127)
$\bigcup_{i=1}^{C}\text{Curvature}(G_{i})$ $\supset\{0,+\kappa,-\kappa,\text{variable}\}$ (128)
View Complementarity Property:
Proposition 1 (Cross-View Information Content).
For any two views $\mathbf{X}^{(h)}$ and $\mathbf{X}^{(h^{\prime})}$ in our synthetic framework:
$\text{MI}(\mathbf{X}^{(h)},\mathbf{X}^{(h^{\prime})})$ $\leq 0.7\cdot\max(\text{H}(\mathbf{X}^{(h)}),\text{H}(\mathbf{X}^{(h^{\prime})}))$ (129)
$\text{MI}(y^{(h)},y^{(h^{\prime})})$ $\geq 0.8\cdot\text{H}(y)$ (130)
ensuring that views provide complementary features while maintaining consistent cluster structure.
6.3.5 Statistical Validation Framework
To ensure the reliability and reproducibility of our synthetic data generation, we implement a comprehensive statistical validation framework:
Shape Fidelity Validation:
For each generated cluster $\mathcal{C}_{k}$, we verify:
Hausdorff Distance: $d_{H}(\mathcal{C}_{k},\mathcal{T}_{k})\leq\epsilon_{shape}$ (131)
Frechet Distance: $d_{F}(\mathcal{C}_{k},\mathcal{T}_{k})\leq\epsilon_{curve}$ (132)
Sample Count Validation: $|\mathcal{C}_{k}|=N_{c}\pm 0$ (133)
where $\mathcal{T}_{k}$ represents the theoretical shape template, $\epsilon_{shape}=0.1$, and $\epsilon_{curve}=0.15$.
Statistical Distribution Testing:
We apply multiple statistical tests to validate the generated data:
  • Kolmogorov-Smirnov test for noise distribution verification
  • Anderson-Darling test for parametric distribution validation
  • Shapiro-Wilk test for Gaussian noise components
  • Chi-square test for uniform angular distributions
All tests are performed with significance level $\alpha=0.05$, and datasets failing any test are regenerated.
@@ 6.4.1 Performance on Synthetic Multi-View Data
Table 2 presents comprehensive performance comparison on our synthetic multi-view dataset. The results demonstrate the superior performance of our proposed heat kernel-enhanced approaches.
@@
Algorithm Accuracy NMI ARI Silhouette
MV-FCM 0.743 0.682 0.651 0.421
CoMSC 0.768 0.701 0.689 0.445
MVDC 0.785 0.734 0.712 0.468
HK-MVFC 1.0000 1.0000 1.0000 1.0000
FedMV-KMeans 0.721 0.663 0.634 0.398
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000
@@
    Significant Performance Gains: HK-MVFC achieves 8-12% improvement in accuracy compared to baseline methods, demonstrating the effectiveness of heat kernel-enhanced distance measures.
@@
    Federated Performance Retention: FedHK-MVFC retains 98.2% of centralized performance while ensuring privacy preservation, indicating minimal performance degradation in federated settings.
@@
    Geometric Structure Capture: The substantial improvement in Silhouette coefficient (40-48% over baselines) indicates superior capability in capturing complex geometric cluster structures.
  4. 4.
    Cross-View Integration: High NMI and ARI scores demonstrate effective integration of complementary information across views.
@@ 6.4.2 Ablation Studies
We evaluate the contribution of different heat-kernel coefficient estimators to clustering performance. The results are summarized in Table 3, comparing the baseline Euclidean distance with two heat-kernel coefficient estimators: Min-Max Normalization (Type 1) and Mean-Variance (Type 2). As can be seen, both heat-kernel methods significantly outperform the baseline, showing the HKC approaches in KED works effectively in enhancing clustering performance.
@@
Estimator Type Accuracy NMI Runtime (s)
@@
For further analysis, we presents the visualization of the heat-kernel coefficient impact. Figure 4 illustrates the differences between the two estimator types. Using Type 1 (Min-Max Normalization) and Type 2 (Mean-Variance) estimators, we observe:
@@
As shown in the figure, the heat-kernel coefficient significantly impacts clustering performance, with Type 1 generally outperforming Type 2 across various metrics. The representation of data through these different scaling methods highlights the importance of feature normalization in clustering tasks. The clustering outcomes demonstrate that the choice of heat-kernel coefficient estimator can lead to substantial differences in clustering quality, with Type 1 providing a more balanced and effective representation of the data. For instance, in scenarios with high feature variability, Type 1’s uniform scaling can mitigate the influence of outliers, resulting in more cohesive clusters. The comprehensive analysis of these effects underscores the necessity of careful feature preprocessing in multi-view clustering, particularly in the context of federated learning where data heterogeneity is prevalent.
@@
$\mathbf{V}_{final}$ $=[0.635,0.365]^{T}\text{ (balanced contribution)}$ (134)
@@
The balanced final weights indicate that both views contribute meaningfully to the clustering solution, validating our complementary data design.
@@
To further understand the clustering behavior, we conduct a misclassification analysis. The confusion matrix reveals specific instances where the HK-MVFC algorithm struggles, particularly in distinguishing between similar phenotypes. By examining these cases, we can identify potential improvements in feature representation and model training. The clustering outcomes is visualized in Figure 6, which shows the distribution of misclassified samples across the four phenotypes.
@@
As shown in the figure 6, the HK-MVFC perfectly distinguishes between the four phenotypes, with no misclassifications observed. This highlights the algorithm’s effectiveness in capturing the underlying structure of the data and its ability to generalize well across different views. The correct predicted labels align perfectly with the ground truth, demonstrating the robustness of the clustering solution.
@@
We also analyze the performance of individual clients in the federated setting. Each client’s contribution to the overall clustering solution is assessed, revealing insights into the data distribution and model performance across different views. This analysis helps identify clients that may require additional support or data augmentation to improve their local models.

### v2

@@ 6.1.1 Mathematical Formulation of View 1
where $\bm{\mu}_{1}=(2,2)^{T}$ is the cluster center, and the radius scaling factor is $0.5$.
@@
where $\bm{\mu}_{2}=(8,2)^{T}$, $a_{2}=1.5$, and $b_{2}=0.4$ are the semi-major and semi-minor axes, respectively.
@@
where $\bm{\mu}_{3}=(2,8)^{T}$ and the horizontal shift of 0.4 in the inner arc create the crescent shape.
@@
6.1.2 Mathematical Formulation of View 2
Data view 2 comprises four distinct patterns that complement those in view 1. These four patterns include the diamond, ring/donut, cross, and heart shapes. The generation of each shape is achieved through the implementation of specific parametric equations, which serve to define their spatial distributions. These complementary information sources, manifested through diverse geometric patterns, are meticulously positioned to generate view-specific clustering challenges. The mathematical formulations employed to generate each cluster in View 2 are as follows:
Cluster 1 (Diamond): Generated using a diamond-shaped polar pattern:
@@ 6.1.2 Mathematical Formulation of View 2
where $\bm{\mu}_{4}^{(2)}=(-2,-2)^{T}$, $\sigma=0.3$ are the scale parameters, and $\eta_{i,x},\eta_{i,y}\sim\mathcal{N}(0,1)$.
@@ 6.1.3 Data Generation Properties and Validation
The synthetic data generation framework ensures several critical properties for rigorous algorithm evaluation. Initially, the issue of sample count consistency must be addressed. It is noteworthy that each cluster in both views contains precisely $N_{c}=2,500$ samples, thereby resulting in a total dataset size of $N=10,000$ samples per view. This balanced design is intended to eliminate any potential bias that may be present in favor of larger clusters.
Secondly, it is imperative to ascertain geometric complexity. The shapes selected for analysis demonstrate a range of geometric complexity, encompassing both convex and non-convex structures. Convex shapes include circles, ellipses, and diamonds, while non-convex shapes encompass crescents, hearts, and crosses. Furthermore, topological variations are observed, including simple connected structures, which constitute the majority of shapes, and multiply connected structures, which include rings. Finally, parametric complexity is considered, with linear structures, such as crosses, and highly nonlinear structures, such as spirals and hearts, being distinguished.
The third component is View Complementarity. The positioning strategy guarantees that:
@@
In this context, the symbol ”MI” is used to represent ”mutual information.” This concept is employed to measure the amount of complementary information provided by views, as opposed to redundant information.
Finally, the Noise Characteristics section provides a comprehensive overview of the subject. Gaussian noise is incorporated with view-specific parameters:
@@ 6.1.4 Federated Data Partitioning Strategy
As delineated in Eqs. (126) and (127), the partitioning scheme guarantees that Client 1 will receive 85% of the data, amounting to 8,500 samples, while Client 2 will receive 15%, or 1,500 samples. It is noteworthy that both clients are equipped with four distinct cluster types. The partition reflects realistic federated scenarios with heterogeneous data distributions. Consequently, this synthetic dataset offers a controlled environment for evaluating multi-view clustering algorithms on geometrically complex data while maintaining known ground truth labels for quantitative performance assessment. As illustrated in Table 1, the synthetic multi-view dataset is characterized by several key statistical properties that render it particularly well-suited for rigorous algorithm evaluation.
@@
The synthetic framework, meticulously designed in Table 1, confers a series of experimental benefits.
@@
In order to guarantee the caliber and dependability of the synthetic dataset, a thorough validation protocol has been executed. This protocol encompasses the following critical assessments:
@@ 6.1.5 Medical Federated Scenario: Multi-Hospital Patient Data Analysis
Initially, Clinical Motivation must be considered. The contemporary healthcare sector produces a substantial amount of multi-modal patient data across various institutions, thereby creating opportunities for collaborative analysis that could enhance patient outcomes through the process of phenotyping and the optimization of treatment. Nevertheless, the existence of stringent privacy regulations (e.g., HIPAA, GDPR) and institutional policies has led to significant impediments to the direct exchange of data between healthcare providers. The FedHK-MVFC framework has been developed to address this challenge by enabling collaborative clustering analysis while ensuring that sensitive patient data never leaves the originating institution.
One plausible scenario that merits consideration is the collaborative identification of cardiovascular patient phenotypes by multiple hospitals, with the objective of enhancing treatment protocols. Each hospital has collected complementary data modalities but lacks the statistical power for comprehensive analysis when working in isolation. By leveraging federated clustering, these institutions can harness the collective knowledge of its constituents while ensuring strict adherence to privacy regulations.
Data Modality Mapping: As demonstrated in Figure 3, the data presented in the form of toy models functions as proxies for actual medical data characteristics. The initial perspective, designated as ”View 1,” pertains to physiological measurements. The four distinct shapes observed in View 1 are indicative of different patient populations based on physiological measurements. The circular cluster thus represents healthy control patients with normal, tightly distributed physiological parameters, while the elliptical cluster models patients with mild cardiovascular risk, showing elongated parameter distributions. The crescent cluster is indicative of patients with complex, non-linear disease progression patterns, while the spiral cluster is representative of patients with progressive disease states demonstrating temporal evolution. The second perspective encompasses both imaging and behavioral data. The complementary shapes in View 2 correspond to imaging and behavioral characteristics. The presence of a diamond cluster is indicative of structured imaging features with clearly delineated anatomical boundaries, while the ring cluster encompasses patients with preserved central function but peripheral abnormalities. The cross cluster reflects behavioral patterns, showing distinct lifestyle factor combinations. Conversely, the presence of the heart cluster signifies the existence of intricate cardiac structural abnormalities, necessitating a specialized interpretation approach.
6.2 Privacy-Preserving Collaborative Analysis
@@
6.2.1 Clinical Workflow Implementation
@@ 6.2.2 Scenario Description
6.2.3 Multi-View Medical Data Structure
@@ 6.2.4 Clustering Objectives
6.2.5 Privacy and Compliance Requirements
@@
6.3 Federated Data Distribution Analysis
@@
6.4 FedHK-MVFC Advantages in Medical Applications
@@
This medical scenario demonstrates how the synthetic data experiments (with their controlled geometric complexities) translate to real-world healthcare applications, where FedHK-MVFC can enable privacy-preserving collaborative analysis across medical institutions.
@@ 6.5 Experimental Setup and Evaluation Metrics
The experimental evaluation of the proposed HK-MVFC and FedHK-MVFC algorithms is conducted using the synthetic multi-view dataset described in the previous subsection. In this study, we will be comparing our proposed HK-MVFC and FedHK-MVFC algorithms against state-of-the-art multi-view clustering methods, both in centralized and federated settings. The centralized multi-view clustering baseline incorporates Co-FKM [cleuziou2009cofkm], MinMax-FCM [wang2017multi], WV-Co-FCM [6862861], and Co-FW-MVFCM [yang2021collaborative]. In the context of federated multi-view clustering, the implementation of Fed-MVFCM [hu2023efficient] is undertaken. The performance of these MVC algorithms in terms of clustering will be evaluated using five external indices: the Adjusted Rand Index (AR), the Rand index (RI) [rand1971objective], the Jaccard index (JI) [jaccard1901distribution], the Fowlkes and Mallows index (FMI) [fowlkes1983method], and the normalized mutual information (NMI) [cover1999elements]. These metrics allow for the quantification of the proportion of misclassified samples relative to the ground truth labels. The experiment is conducted using Matlab R2025a on a workstation equipped with an Intel Core i9-12900K central processing unit (CPU), 32 gigabytes of random access memory (RAM), and an NVIDIA RTX 3090 graphics processing unit (GPU). Each algorithm is executed ten times with different random initializations to ensure robustness, and the average performance metrics are reported.
@@ 6.6.1 Performance on Synthetic Multi-View Data
Table 2 presents a comprehensive performance comparison of our synthetic multi-view dataset. The results demonstrate the superior performance of our proposed heat kernel-enhanced approaches.
@@
Algorithm AR NMI RI JI FMI Runtime (s)
Centralized MVC
Co-FKM 0.4391 0.3979 0.6484 0.3591 0.5466 12.3
MinMax-FCM 0.9776 0.9881 0.9906 0.9750 0.9848 36.03
WV-Co-FCM 0.2500 0.0000 0.2499 0.2499 0.4999 256.2
Co-FW-MVFCM 0.7835 0.8270 0.9165 0.7442 0.8394 122.5
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000 15.7
Federated MVC
Fed-MVFCM 0.9926 0.9723 0.9927 0.9711 0.9853 136.828
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000 28.4
@@
    The Superior Ones: As reported in Table 2, the proposed algorithms HK-MVFC and FedHK-MVFC consistently outperform Co-FKM, MinMax-FCM, WV-Co-FCM, and Co-FW-MVFCM with respect to the average values of AR, NMI, RI, JI, and FMI. The only exception is MinMax-FCM for NMI and JI, which also exhibits superior performance compared with Fed-MVFCM. Overall, these results indicate that HK-MVFC and FedHK-MVFC achieve the most favorable comprehensive performance among all evaluated algorithms. The high AR, RI, JI, NMI, and FMI scores further demonstrate that the QFT-based approach effectively integrates complementary information across multiple views.
@@
    Significant Performance Gains: HK-MVFC attains a 12–50% increase in clustering accuracy relative to centralized approaches, including Co-FKM, WV-Co-FCM, and Co-FW-MVFCM, and achieves a 1–3% improvement across all five evaluation metrics compared to the decentralized method Fed-MVFCM. These results substantiate the effectiveness of incorporating heat kernel–enhanced distance measures into the clustering framework.
@@
    Federated Performance Retention: FedHK-MVFC maintains a level of accuracy comparable to that of the centralized setting while preserving privacy guarantees, thus exhibiting negligible performance degradation under federated learning conditions. Its runtime is approximately twice that of its baseline method, HK-MVFC, which is anticipated given that FedHK-MVFC introduces multiple additional computational phases and server-side operations. Compared to its predecessor, Fed-MVFCM, FedHK-MVFC executes more efficiently. Specifically, FedHK-MVFC achieves convergence earlier and requires fewer communication rounds (aggregation stages) than Fed-MVFCM (see Table 9).
@@ 6.6.2 Ablation Studies
For the centralized algorithms Co-FKM, MinMax-FCM, WV-Co-FCM, and Co-FW-MVFCM, as well as for the decentralized Fed-MVFCM, we varied the fuzzifier parameter over the set $m\in\{1.05,1.25,1.50,1.75,2.00\}$ while keeping all other hyperparameters fixed. The corresponding results are reported in Tables 37. These ablation experiments indicate that the proposed HK-MVFC and FedHK-MVFC methods consistently outperform state-of-the-art centralized and decentralized MVFCM approaches.
Table 3: Performance Comparison on Synthetic Multi-View Data in Terms of Average Adjusted Rand Index (AvG-AR) Under Varying Fuzzifier Values $m$
Algorithm $m=1.05$ $m=1.25$ $m=1.50$ $m=1.75$ $m=2.00$
Centralized MVC
Co-FKM 0.5000 0.5000 0.5045 0.5411 0.4200
MinMax-FCM 0.9254 0.9625 1.0000 1.0000 1.0000
WV-Co-FCM 0.2500 0.2500 0.2500 0.2500 0.2500
Co-FW-MVFCM 0.5338 0.8510 0.8685 0.8354 0.8247
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Federated MVC
Fed-MVFCM 0.6463 0.7597 0.9919 0.9926 0.9926
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Table 4: Performance Comparison on Synthetic Multi-View Data in Terms of Average Rand Index (AvG-RI) Under Varying Fuzzifier Values $m$
Algorithm $m=1.05$ $m=1.25$ $m=1.50$ $m=1.75$ $m=2.00$
Centralized MVC
Co-FKM 0.6704 0.6704 0.6000 0.6582 0.6429
MinMax-FCM 0.9688 0.9844 1.0000 1.0000 1.0000
WV-Co-FCM 0.2499 0.2499 0.2499 0.2499 0.2499
Co-FW-MVFCM 0..8227 0.9423 0.9490 0.9363 0.9321
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Federated MVC
Fed-MVFCM 0.6815 0.8483 0.9919 0.9926 0.9927
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Table 5: Performance Comparison on Synthetic Multi-View Data in Terms of Average Normalized Mutual Information (AvG-NMI) Under Varying Fuzzifier Values $m$
Algorithm $m=1.05$ $m=1.25$ $m=1.50$ $m=1.75$ $m=2.00$
Centralized MVC
Co-FKM 0.4510 0.4510 0.3663 0.4233 0.2980
MinMax-FCM 0.9688 0.9802 1.0000 1.0000 1.0000
WV-Co-FCM 0.0000 0.0000 0.0000 0.0000 0.0000
Co-FW-MVFCM 0.6392 0.8764 0.8900 0.8674 0.8621
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Federated MVC
Fed-MVFCM 0.5634 0.7898 0.9696 0.9722 0.9723
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Table 6: Performance Comparison on Synthetic Multi-View Data in Terms of Average Fowlkes and Mallows Index (AvG-FMI) Under Varying Fuzzifier Values $m$
Algorithm $m=1.05$ $m=1.25$ $m=1.50$ $m=1.75$ $m=2.00$
Centralized MVC
Co-FKM 0.5665 0.5665 0.5546 0.5686 0.4766
MinMax-FCM 0.9492 0.9746 1.0000 1.0000 1.0000
WV-Co-FCM 0.4999 0.4999 0.4999 0.4999 0.4999
Co-FW-MVFCM 0.6532 0.8905 0.9035 0.9789 0.8711
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Federated MVC
Fed-MVFCM 0.7594 0.8255 0.9839 0.9853 0.9854
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Table 7: Performance Comparison on Synthetic Multi-View Data in Terms of Average Jaccard Index (AvG-JI) Under Varying Fuzzifier Values $m$
Algorithm $m=1.05$ $m=1.25$ $m=1.50$ $m=1.75$ $m=2.00$
Centralized MVC
Co-FKM 0.3764 0.3764 0.3562 0.3788 0.3075
MinMax-FCM 0.9167 0.9583 1.0000 1.0000 1.0000
WV-Co-FCM 0.2499 0.2499 0.2499 0.2499 0.2499
Co-FW-MVFCM 0.5188 0.8078 0.8301 0.7887 0.7758
HK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
Federated MVC
Fed-MVFCM 0.6338 0.7122 0.9683 0.9710 0.9711
FedHK-MVFC 1.0000 1.0000 1.0000 1.0000 1.0000
@@
We evaluate the contribution of different heat-kernel coefficient estimators to clustering performance. The results are summarized in Table 8, comparing the baseline Euclidean distance with two heat-kernel coefficient estimators: Min-Max Normalization (Type 1) and Mean-Variance (Type 2). As can be seen, both heat-kernel methods significantly outperform the baseline, indicating that the HKC approaches in KED work effectively to enhance clustering performance.
@@
Estimator Type AR NMI Runtime (s)
@@
For further analysis, we present the visualization of the impact of the heat-kernel coefficient. Figure 6 illustrates the differences between the two types of estimators. Using Type 1 (Min-Max Normalization) and Type 2 (Mean-Variance) estimators, we observe:
@@
As shown in the figure, the heat-kernel coefficient significantly impacts clustering performance, with Type 1 generally outperforming Type 2 across various metrics. The representation of data through these different scaling methods highlights the importance of feature normalization in clustering tasks. The clustering outcomes demonstrate that the choice of heat-kernel coefficient estimator can lead to substantial differences in clustering quality, with Type 1 providing a more balanced and effective representation of the data. For instance, in scenarios with high feature variability, Type 1’s uniform scaling can mitigate the influence of outliers, resulting in more cohesive clusters. The comprehensive analysis of these effects underscores the necessity of careful feature preprocessing in multi-view clustering, particularly in the context of federated learning, where data heterogeneity is prevalent.
@@
$\mathbf{V}_{final}$ $=[0.635,0.365]^{T}\text{ (imbalanced contribution)}$ (133)
@@
The non-uniform distribution of the final weights indicates that the two views make differential contributions to the clustering solution, thereby corroborating the validity of our complementary data design. In the distributed setting reported in Figures 9-10, the final local stage for each client revealed that, for two clients, data view 1 contributed more substantially than data view 2 to the derived clustering solution (see Figure 11).
@@
To further understand the clustering behavior, we conduct a misclassification analysis. The confusion matrix reveals specific instances where the HK-MVFC algorithm struggles, particularly in distinguishing between similar phenotypes. By examining these cases, we can identify potential improvements in feature representation and model training. The clustering outcomes are visualized in Figure 8, which shows the distribution of misclassified samples across the four phenotypes.
@@
As demonstrated in Figure 8, the HK-MVFC model exhibits an optimal capacity for distinguishing between the four distinct phenotypes, with no instances of misclassification observed. This underscores the efficacy of the algorithm in capturing the underlying structure of the data and its capacity to generalize effectively across diverse perspectives. The predicted labels that were deemed correct align perfectly with the ground truth, thereby demonstrating the robustness of the clustering solution.
@@
We also analyze the performance of individual clients in the federated setting. Each client’s contribution to the overall clustering solution is assessed, revealing insights into data distribution and model performance across different views. This analysis helps identify clients who may require additional support or data augmentation to improve their local models.